40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 10 11:55:22 2021, gautam, Update, SUS, PRM suspension suspect PRM_sensorSpectra.pdfPRM_pkFitNaive.pdfPRM_diagComp.pdf
    Reply  Wed Mar 10 17:35:47 2021, gautam, Update, SUS, Spooky action at a distance PRMtrip.pngPRMtrip_zoom.png
Message ID: 15894     Entry time: Wed Mar 10 11:55:22 2021     Reply to this: 15898
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRM suspension suspect 

The procedure is that the optic is kicked to excite it, and allowed to ring down for ~1ksec, with damping turned off. The procedure is repeated 15 times for some averaging. 

Attachment #1 - sensor spectra from yesterday.

Attachment #2 - peaks using the naive diagonalization matrix from yesterday.

Attachment #3 - Data from ~1 year ago. 

The y-axis in all plots is labelled as "cts/rtHz" but these are the DQed channels, which come after a "cts2um" CDS filter - so if that filter is accurate, them the y-axes may be read as um/rtHz.

I wonder if the September 2020 earthquake somehow damaged the PRM suspension, as this experiment would suggest that the problem is not only with the actuation. The data was gathered with the neutral position of the PRM (between kicks) being well aligned for PRMI, and the DC values of all the shadow sensors in this position is close to half-light (~1V, except for side which was more like 4V). Hard to say what exactly is happening since only the PIT DoF has the weird asymmetric peak shape instead of the expected Lorentzian - I would have thought that a damaged wire or broken magnet would affect all 4 DoFs but the F.C. spring experience on ETMY showed that anything is possible.

Attachment 1: PRM_sensorSpectra.pdf  338 kB  | Hide | Hide all
PRM_sensorSpectra.pdf
Attachment 2: PRM_pkFitNaive.pdf  440 kB  | Hide | Hide all
PRM_pkFitNaive.pdf
Attachment 3: PRM_diagComp.pdf  67 kB  | Hide | Hide all
PRM_diagComp.pdf
ELOG V3.1.3-